Fermi National Laboratory


Accelerator Update for July 6, 2005 - July 8, 2005

Wednesday July 6
The day shift began with Operations monitoring store 4250 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the following experiments taking beam: the Collider Detector at Fermilab (CDF), D-Zero (D0), MiniBooNE, NuMI, and MCenter.

Operations terminated store 4250 at 9:01 AM.

Operators prepared CDF and D0 for an access into their collision halls at 9:14 AM. Also, a Booster expert began a study period.

CDF came out of access at 10:03 AM.

Operations reported at 10:17 AM that cryo system technicians were conducting maintenance on wet engines for Tevatron (TeV) sectors D1 and E3.

Operations reported that experts from the Main Injector (MI) and Pbar were conducting studies.

D0 came out of access at 11 AM.

The Cryo Coordinator reported at 11:02 AM that they had completed their maintenance work.

Operations put the TeV through a dry squeeze at 12:40 PM.

Operations put the TeV into shot setup at 1:03 PM.

At 1:51 PM, the Cryo Coordinator reported a problem with the TeV sector D1 cold compressor. This put shot setup on hold while techs investigated.

Operations established combination store 4252 at 4:59 PM with an initial luminosity of 104.67E30.

At 5:13 PM, Operations reported that stacking was on hold due to an Accumulator SEPTA problem (A:ISEP2V). An expert investigated. Pbar resumed stacking at 5:47 PM.

Operations established beam to the Meson MCenter experiment at 7:46 PM.

A Pbar expert began a stacking study at 8:05 PM.

Operations put the Recycler into shot setup at 9:47 PM to stash antiprotons. The transfers finished some time around 10:39 PM. At least one of the transfers didnít make it from Pbar to the Recycler.

Pbar resumed stacking at 11:11 PM. Operations also reported problems with the MI RF system; stations 1, 3, 4, 5, and 7 all tripped off. When the MI RF system trips off, NuMI also goes down. Operations contacted an expert.

At 11:54 PM, Operations reported that they had rebooted the MI low level RF, but this made things worse. Operations called in experts.

Thursday July 7
The midnight shift began with Operations monitoring store 4252 and the stash, with Pbar off due to a SEPTA problem, and with the following experiments taking beam: CDF, D0, and MiniBooNE.

At 12:14 AM, Operations reported that A:ISEP2V tripped off again. Operators investigated.

At 1:26 AM, Operations reported MI damper problems.

A damper expert re-initialized the MI dampers around 2:23 AM, which allowed Operations to run a NuMI-only Time Line Generator (TLG) module. Operators replaced a ramp card for a power supply (I:F17B3) and then called a power supply expert. This power supply was tripping off the Pbar beam permit.

The power supply expert instructed an operator to adjust the flattop value for I:F17B3. Pbar resumed stacking at 2:41 AM.

Operations reported at 5:26 AM that experts had returned the dampers to normal operation, but that there were still MI problems.

At 6:05 AM, another MI expert reported that she found a beam loading problem. She and another expert contacted a third and all three investigated further. Stacking and NuMI beam resumed normal operations at 7:08 AM.

At 7:55 AM, Operations reported that there was still MI RF problems and that experts were out at MI-60 investigating.

Operations reported at 10:42 AM that techs had completed their work on LAM62.

Operations reported at 10:59 AM that techs had completed their work on ISEP2V.

A Linac technician reported at 11:12 AM that he had finished work on LRF1 and was nearly finished with LRF4,

RF experts reported at 11:28 AM that MIRF10 was back on line.

Operations had established beam to Pbar and NuMI by 12:30 PM.

Operations established beam to MCenter at 12:56 PM.

Pbar experts began a reverse proton study at 1:35 PM.

The Pbar study ended and Pbar resumed stacking a 5:04 PM.

At 5:56 PM, Operations reported that LRF4 was down due to bad vacuum in its power amplifier. An operator called an expert who suggested that operators adjust the current.

MiniBooNE experimenters requested their horn to be turned off at 6:01 PM.

At 8:46 PM, LRF4 tripped off due to a bad vacuum for the first of many times over the next two hours. There were a couple of twenty to thirty minute conditioning periods in that time frame. When LRF4 was off, there was no downstream beam.

A MI expert reported at 9:20 PM that our MI problems were due to the feed forward gain levels during slipstacking.

Operations terminated store 4252 at 11:05 PM.

Operations had D0 prepared for access by 11:24 PM.

Friday July 8
The midnight shift began with Operations monitoring the stash and stack, with D0 in access, and with all beam down due to trouble with LRF4.

Operations reported that LRF4 came back on at 12:17 AM. Beam was reestablished.

D0 came out of access at 1:02 AM.

At 1:16 AM, LRF4 tripped off. Operators had it back up two minutes later.

Operations put the TeV into a wet squeeze at 1:42 AM.

Operations put the TeV into shot setup at 2:07 AM.

At 3:01 AM, LRF4 tripped off. Operators had it back up about two minutes later.

At 3:23 AM, Operations reported MI coalescing problems. Operations contacted an expert.

At 3:27 AM, LRF4 tripped off again.

LRF4 tripped off twice more at 4:37 AM. It reset okay.

Operations prepared NuMI for access at 5:17 AM.

Operations established combination store 4254 at 5:37 AM with an initial luminosity of 81.3E30.

Pbar resumed stacking 5:35 AM.

MCenter resumed taking beam at 5:52 AM.

At 6:30 AM, Operations reported that they were not able to keep the MI beam intensity up at the levels needed.

At 6:38 AM, LRF4 tripped off. It reset okay.

At 6:55 AM, Operations reported stacking problems due to numerous Debuncher trim magnet trips. Operators investigated and then called a Pbar expert and called in controls experts.

At 7:04 AM, LRF4 tripped off. It took operators nine attempts to get it back on, and then it tripped off again at 7:24 AM.

LRF4 came back on at 7:37 AM, but then KRF5 tripped off. When KRF5 came back on LRF4 tripped off. A Linac expert investigated.

Operations reported at 8:05 AM, that the Debuncher trim magnets resumed functioning normally while the controls experts were investigating.

The Plan for Friday and the Weekend
The plans for today are to allow some much needed MI, Pbar, and Recycler studies. The plans for the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/5/2005   email Fermilab